OpenStudy (anonymous):

A 9 foot ladder is 3 feet from the base of the wall. how high is the ladder on the wall ?round answer to the nearest 1000

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

a=3; c=9

OpenStudy (anonymous):

3^2+b^2=9^2

OpenStudy (anonymous):

sqrt(81-9)=b

OpenStudy (anonymous):

Round to the nearest thousand!? Are you sure about that @dmmeke ?